AIR-BROADENED LINE SHAPES IN THE 2ν_{3} R BRANCH OF ^{12}CH_{4} BETWEEN 6014 AND 6100 CM^{−1}

\underline{V. MALATHY DEVI}, D. CHRIS BENNER, Dept. of Physics, The College of William and Mary, Williamsburg, VA 23187; K. SUNG, L. R. BROWN, T. J. CRAWFORD, SHANSHAN YU, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109; M. A. H. SMITH, Science Directorate, NASA Langley Research Center, Hampton, VA 23681; A. W. MANTZ, Dept. of Physics, Astronomy and Geophysics, Connecticut College, New London, CT 06320.

Complete and accurate information on line shape parameters of 2ν_{3} methane transitions for air broadening as a function of temperature is critical not only for the correct interpretation of the observed atmospheric spectra but also for the development of a reliable theoretical model. For this reason, we obtained a series of high-resolution, high S/N spectra of high-purity ^{12}CH_{4} and ^{12}CH_{4} broadened with dry air at temperatures in the 130 to 295 K range using the Bruker IFS 125HR Fourier transform spectrometer at JPL. Two absorption cells were used in the experiment, a White cell with path length of 13 m for room temperature spectra and a 21 m Herriott cell^{a} for cold sample spectra. The 15 spectra used in the analysis consisted of 3 low pressure (0.26 to 2.57 Torr) spectra with pure ^{12}CH_{4} and 12 air-broadened spectra with total sample pressures of 79-805 Torr and volume mixing ratios of methane between 0.23 and 1%. A multispectrum least-squares fitting technique^{b} was employed to fit all 15 spectra simultaneously. Preliminary results for select R(J) manifolds will be presented.^{c}

aA. W. Mantz, K. Sung, L. R. Brown, et al., abstract submitted to this Symposium.

bD. C. Benner, C. P. Rinsland, V. Malathy Devi, M. A. H. Smith and D. Atkins, JQSRT \underline{53} (1995) 705-721.

cResearch described in this paper was performed at Connecticut College, the College of William and Mary, NASA Langley Research Center and the Jet Propulsion Laboratory, California Institute of Technology, under contracts and cooperative agreements with the National Aeronautics and Space Administration.